OpenStudy (anonymous):

1. Julia deposited $250 in a savings account that earns 2.7% simple interest. How much interest has Julia earned by the end of the first year? $6.75 $92.59 $256.75 $675.00 2. Armand deposited $389.42 in a savings account that earns 3.2% simple interest. What is Armand’s account balance after seven years? $87.23 $401.88 $476.65 $872.30 3. Andy deposited $1,567.12 in a savings account that earns 1.9% simple interest. What will Andy’s account balance be in nine months? $1,567.12 $1,589.45 $1,596.90 $1,835.10

OpenStudy (whpalmer4):

Interest earned with simple interest is just PV * i * t where PV is present value (the balance at the start of the problem), i is the interest rate expressed as a decimal (so divide the percentage by 100, 2.3% = 2.3/100 = 0.023), and t is the number of years (be sure to convert to years if you get something else). To find a future balance (FV), FV = PV + PV*i*t 1. PV = $250, i = 2.7% = 0.027, t = 1 interest earned = PV*i*t = $250*0.027*1 = $6.75 Easy, right? 2. PV = $389.42, i = 3.2% = 0.032, t = 7 Here we want to find the FV, not the interest earned FV = PV + PV*i*t = $389.42 + $389.42*0.032*t = 3. PV = $1,567.12, i = 1.9% = 0.019, t = 9 months = 9/12 = 0.75 Note that we had to convert months into years — the answer would be very incorrect if you did not do so! Here we are finding the FV FV = PV + PV*i*t = $1,567.12 + $1,567.12*0.019*0.75 = You have to read these problems carefully to make sure you find the right quantity!